# soft open fences

### v2

@@ 1 Introduction
  • We reveal that EM is a distribution shaping tool rather than a learning method by analyzing the inconsistency in loss-reasoning curve 3.4 and the logit shift 3.3 effect of EM.
@@ 3.1 Experimental Setting
Temperature Batch Size Learning Rate Max Response Len
0.5 64 $2\times 10^{-5}$ 2048
Table 1: Training hyper-parameters
We implemented the overall training process of EM based on Acclerate [7]. We selected 1 piece of data from the dataset as prompt. Since it is an unsupervised method, we do not need any data labels. We directly train the model for only 10 steps with a constant learning rate of $2\times 10^{-5}$, a temperature parameter of 0.5, and a batch size of 64.  [^1] The reason why only 10 steps are sufficient will be explained in detail in Section 3.4.
Model Dataset Size Training Step MATH500 Minerva Math Olympiad Bench AMC23 Avg.
OpenReasoner-Zero-7B[9] 129k 600+ 79.2 31.6 44.0 47.0 50.5
SimpleRL-Zoo[24] 24K 4000 76.8 30.9 39.4 55.3 50.6
Prime-Zero-7B[3] 230K 240 83.8 36.0 40.9 62.7 55.9
Oat-Zero-7B[12] 12K 300 80.0 30.1 41.0 62.7 53.5
RLVR-GRPO[20] 1.2 K 1000 78.6 33.8 41.6 62.5 54.1
+ 16-shot 16 1000 77.8 35.3 39.9 62.2 53.8
+ 1-shot 1 1000 78.6 36.0 43.7 61.9 55.1
Qwen2.5-Math-7B NA NA 53.0 11.0 17.2 44.1 31.3
+ EM 1-shot 1 10 78.8↑25.8 35.3↑24.3 39.7↑22.5 70.3↑26.2 56.0↑24.7
Table 2: Comparison of different methods on math reasoning benchmarks.
@@ 3.3 Logits Shift
As shown in Figure 2, the loss drops to a relatively low level around step 10, and the model’s performance on mathematical reasoning reaches its peak. However, unexpectedly, as the EM training loss continues to decrease beyond step 10, the mathematical reasoning performance begins to decline. In conjunction with the findings presented in Section 3.3, we argue that EM functions primarily as a tool for shaping the model’s distribution rather than as a learning method or strategy. Consequently, the effect of distribution shaping is largely achieved within a very small number of training steps, leading to a decoupling between the continued decrease in EM training loss and improvements in mathematical reasoning performance.
@@ 3.6 Sampling Temperature in Evaluate Generation
Together with our analysis in Section 3.3, we hypothesize that EM training systematically reshapes the model’s logits distributions to become increasingly right-skewed. This process reinforces confidence in already high-probability tokens, effectively concentrating the probability mass on semantically coherent and correct options. As a result, greedy decoding—which deterministically selects the most probable token—becomes a highly effective strategy post-EM fine-tuning.
@@ 3.8 EM on Various Models
Model MATH500 Minerva Math Olympiad Bench AMC23 Avg.
LLaMA-3.1-8B 49.0 21.3 15.4 25.6 27.8
+ EM multi-shot 49.2 23.2 13.8 25.9 28.0
+ EM 1-shot 42.6 21 17 37.8 29.6
Qwen2.5-7B 59.8 10.3 29.2 40.3 34.9
+ EM multi-shot 61.8 18.4 31.1 43.4 38.7
+ EM 1-shot 67.4 22.1 33.6 45.6 42.2
Qwen2.5-7B-Instruct 77.0 37.1 37.9 51.9 50.9
+ EM multi-shot 74.6 34.2 36.4 48.4 48.4
+ EM 1-shot 76.4 36.4 39.3 54.1 51.5
SimpleRL-Zoo 76.8 30.9 39.4 55.3 50.6
+ EM multi-shot 77.2 29.0 35.7 55.9 49.5
+ EM 1-shot 73.0 24.6 35.3 45.9 44.7
Qwen2.5-Math-7B 53.0 11.0 17.2 44.1 31.3
+ EM multi-shot 68.0 18.8 33.2 61.6 45.4
+ EM 1-shot 78.8 35.3 39.7 70.3 56.0
Table 3: Comparison of different models on various mathematical benchmarks.
@@ 5.1 Entropy Minimization
Agarwal et al.[20] was the first to discover that simply optimizing the entropy loss can lead to significant improvements in reasoning performance. However, they did not explore this phenomenon in depth, stating in the original paper, "We leave the rigorous analysis to future works." Wang et al.[1] is the first work to study entropy minimization in post-training for large language models. However, they consider the effectiveness of entropy minimization to be unreasonable, presenting experimental results but offering limited in-depth analysis of entropy minimization.

### v3

@@ 1 Introduction
  • We reveal that EM is a distribution shaping tool rather than a learning method by analyzing the inconsistency in loss-reasoning curve over confidence in Section 3.4) and the logit shift(in Section 3.3) effect of EM.
@@ 3.1 Experimental Setting
We implemented the overall training process of EM based on Acclerate [8]. We selected 1 piece of data from the dataset as prompt. Since it is an unsupervised method, we do not need any data labels. We directly train the model for only 10 steps with a constant learning rate of $2\times 10^{-5}$, a temperature of 0.5, and a batch size of 64. The reason why only 10 steps are sufficient will be explained in detail in Section 3.4.
Model Dataset Size Training Step MATH 500 Minerva Math Olympiad Bench AMC23 KK MBPP Avg.
OpenReasoner-Zero-7B[11] 129k 600+ 79.2 31.6 44.0 47.0 27.2 78.3 51.1
SimpleRL-Zoo[28] 24K 4000 76.8 30.9 39.4 55.3 17.4 75.9 49.3
Prime-Zero-7B[4] 230K 240 83.8 36.0 40.9 62.7 4.4 41.0 44.8
Oat-Zero-7B[15] 12K 300 80.0 30.1 41.0 62.7 2.2 58.7 45.8
RLVR[23] 1.2 K 1000 78.6 33.8 41.6 62.5 12.4 54.5 47.2
RLVR 16-shot 16 1000 77.8 35.3 39.9 62.2 2.4 58.5 46.0
RLVR 1-shot 1 1000 78.6 36.0 43.7 61.9 5.8 60.3 47.7
Qwen2.5-Math-7B NA NA 53.0 11.0 17.2 44.1 1.0 48.9 29.2
+ EM 1-shot 1 10 78.8↑25.8 35.3↑24.3 39.7↑22.5 70.3↑26.2 17.4↑16.4 65.1↑16.2 51.1↑24.9
Table 1: Comparison of different methods on math reasoning benchmarks (MATH500[14], MinervaMath[12], OlympiadBench[9], AMC23), logic reasoning benchmark (KK[24]), and code benchmark (MBPP[2]). To reduce randomness, each benchmark is evaluated under avg@8.
@@ 3.3 Logits Shift
As shown in Figure 2, the loss drops to a relatively low level around step 10, and the model’s performance on mathematical reasoning reaches its peak. However, unexpectedly, as the EM training loss continues to decrease beyond step 10, the mathematical reasoning performance begins to decline. We designate this phenomenon as over confidence. Persistent EM may excessively amplify the model’s confidence in its tokens during inference, thereby exacerbating algorithmic bias and leading to significant deviations in outputs. In conjunction with the findings presented in Section 3.3, we argue that EM functions primarily as a tool for shaping the model’s distribution rather than as a learning method or strategy. Consequently, the effect of distribution shaping is largely achieved within a very small number of training steps, leading to a decoupling between the continued decrease in EM training loss and improvements in mathematical reasoning performance.
@@ 3.6 Sampling Temperature in Evaluate Generation
Together with our analysis in Section 3.3, we hypothesize that EM training systematically reshapes the model’s logits distributions to become increasingly right-skewed. This process reinforces confidence in already high-probability tokens, effectively concentrating the probability mass on semantically coherent and correct options. As a result, greedy decoding—which deterministically selects the most probable token—becomes a highly effective strategy after EM training.
@@ 3.8 EM on Various Models
Model MATH 500 Minerva Math Olympiad Bench AMC23 KK MBPP Avg.
LLaMA-3.1-8B 49.0 21.3 15.4 25.6 7.6 73.0 32.0
+ EM multi-shot 49.2 23.2 13.8 25.9 7.6 73.0 32.1
+ EM 1-shot 42.6 21.0 17.0 37.8 10.2 70.1 33.1
Qwen2.5-7B 59.8 10.3 29.2 40.3 11.8 72.5 37.3
+ EM multi-shot 61.8 18.4 31.1 43.4 14.0 78.3 41.2
+ EM 1-shot 67.4 22.1 33.6 45.6 12.6 75.9 42.9
Qwen2.5-7B-Instruct 77.0 37.1 37.9 51.9 17.0 80.7 50.3
+ EM multi-shot 74.6 34.2 36.4 48.4 11.0 78.3 47.2
+ EM 1-shot 76.4 36.4 39.3 54.1 14.8 77.5 49.8
SimpleRL-Zoo 76.8 30.9 39.4 55.3 17.4 75.9 49.3
+ EM multi-shot 77.2 29.0 35.7 55.9 17.8 73.8 48.2
+ EM 1-shot 73.0 24.6 35.3 45.9 15.4 73.0 44.5
Qwen2.5-Math-7B 53.0 11.0 17.2 44.1 1.0 48.9 29.2
+ EM multi-shot 68.0 18.8 33.2 61.6 9.0 67.2 43.0
+ EM 1-shot 78.8 35.3 39.7 70.3 17.4 65.1 51.1
Table 2: Comparison of different methods on math reasoning benchmarks (MATH500[14], MinervaMath[12], OlympiadBench[9], AMC23), logic reasoning benchmark (KK[24]), and code benchmark (MBPP[2]). To reduce randomness, each benchmark is evaluated under avg@8.
@@ 5.1 Entropy Minimization
Wang et al. [23] discovered that RLVR can achieve performance comparable to using thousands of data with just a single data, which is one of the main inspirations for our work. They also were the first to observe that simply optimizing the entropy loss can significantly improve reasoning performance. However, they did not explore this phenomenon in depth, stating in the original paper, “We leave the rigorous analysis to future works.” Agarwal et al. [1] is the first to study entropy minimization during post-training for large language models. However, they considered the effectiveness of entropy minimization to be unreasonable; although they presented experimental results, their analysis of entropy minimization remained limited.